OpenStudy (anonymous):

9. Justify the following statement using an algebraic proof: In a triangle, if the sum of the measures of two angles is equal to the measure of the third angle, then the triangle is a right triangle. statements? reasons?

OpenStudy (anonymous):

let 3 angles be x,y,z x + y = z so: x + y - z = 0 x + y + z = 180 subtract -2z = -180 z = 90 so triangle is right-angled

OpenStudy (anonymous):

do u follow that ok?

OpenStudy (anonymous):

i think i can figure it out a little better now, thank you. the only thing that confuses me is its one of those chart things that have statements/reasons so i need to write down the reasons

OpenStudy (anonymous):

ok well reason for x+y+z = 180 is 3 angles of triangle add up to 180 degrees - excuse me if u knew that already!

OpenStudy (anonymous):

yeah its alright though! it helped a lot!

OpenStudy (anonymous):

good